OpenStudy (anonymous):

A television and DVD player cost a total of $1240 . The cost of the television is three times the cost of the DVD player. Find the cost of each item.

Nnesha (nnesha):

let television = x and DVD = y television and dVD cost 1240 so first equation would be x+y=1240 try to write 2nd equation `The cost of the television is three times the cost of the DVD player` `is` means `=` equal sign

OpenStudy (secret66):

1240 / 3 = 413.3 DVD player = $413.30 1240 - 413.3 = 826.7 Television = $826.70

OpenStudy (shadowlegendx):

^ direct answers, reported